Find the number of moles in 452L of NH3 at STP. 

Find the Volume in 7.86 mol of O2 at STP. 

Calculate the number of Liters in 1483g of NH3 at STP.

How many moles are in 288 L of CO2?

How many grams are in 1288 L of CO2?

Find the mass in grams of 32L of NH3 at STP


1
Expert's answer
2021-03-12T06:02:14-0500

I. V= 425L

Molar volume at STP =22.4L

Avogadro's number= 6.02 x 1023

Mole= volume/molar volume= number of particles/Avogadro's number

Number of particles= Avogadro's number x volume/molar volume

= 6.02x1023 x 425/22.4

= 1.14 x 1025


II. Volume= mole x molar volume

= 7.86 x 22.4= 176L


III. Mass= 1483g

Molar mass of NH3= 17g/mol

Mole= mass/molar mass= volume/Molar volume

Volume= mass x molar volume/molar mass

= 1483x22.4/17= 1954L


IV. Mole= volume/molar volume= 288/22.4= 12.9mol


V. Mole = 1288/22.4= 57.5mol

Mass= mole x Molar mass

= 57.5 x 44= 2530g


VI. Mole= 32/22.4= 1.43mol

Mass= mole x molar mass

= 1.43 x 17= 24.3g